OpenStudy (anonymous):

11. Divide (11x^2 + 2x^3 + 14 + 17x) ÷ (2x + 5).

OpenStudy (anonymous):

first step: order the dividend (first polynomial ) in order of decreasing powers of x

OpenStudy (anonymous):

second step: choose synthetic division or polynomial division

OpenStudy (anonymous):

I dont know which to do ;;

OpenStudy (anonymous):

haha. it depends on your preference. I can go through both if you'd like, but it'll take a while (these things are hard to write out without speech)

OpenStudy (anonymous):

Synthetic Division: solve the divisor 2x+5 = 0 for x 2x = -5 x = -5/2 = -2.5 put -2.5 on the far upper left as seen below. Place the coefficients of the terms of the divisor in the box-thing (see below) in order of decreasing powers of x.  bring the 2 down. multiply this 2 by -2.5 and place it (-5)under the 11 Add -5 to11 to get 6. Multiply this 6 by -2.5 and place it (-15) under 17 Add -15 to 17 to get 2. Multiply this by -2.5 and place it (-5) under 14 Add -5 to 14 to get 9, your remainder

OpenStudy (anonymous):

 so answer is 2x^2 + 6x + 2 remainder 9 or 2x^2 + 6x + 2 + 9/(2x^3 + 11x^2 + 17x + 14)

OpenStudy (anonymous):

Thank you. Sorry for my late replies, lag.